The weight of a can of soup is jointly proportional to its height and to the square of its diameter.
A can 8 inches high with a diameter of 4 inches weighs 51.2 ounces. What is the weight of a can that is 4 inches high with a diameter of 3 ​inches?

Answer by amalm06(224)

W=kh*(d^2)

51.2=(k)(8)(16), so k=0.4

For h=4 and d=3, W=0.4*4*9=14.4 oz (Answer)